SOAR Winter Storm

Submitted by jelias on Fri, 2015-08-14 11:48

(Final update, August 18) Over the last 10 days, SOAR was affected by a severe winter storm that occurred on the weekend of August 7. High winds and rain developed during the afternoon and evening of August 7, at which point the dome was evacuated. All staff were evacuated from the mountain 2 days later, with continuing snowfall and wind. The process of re-opening the road had already begun, but the work to re-open the section from the dormitory/kitchen area to the telescopes (Gemini, SOAR and LSST construction site) proved very slow due to the accumulation of snow – estimated at over a meter on average, and significantly deeper on sections of the road where the wind had pushed it into drifts. Contact with the SOAR computers was lost during the evening of the 8th when the generator stopped operation; this was because snow was blocking the generator’s air intake.

The NOAO operations crew re-opened the road for access to SOAR and the other telescopes on the 13th, and a crew led by Gerardo Gomez went up to restore operations. Instruments were brought on-line in the order they were scheduled for use, although without internet connections off the summit (this process continues through August 18). Fortunately, the observers are on-site so the absence of the internet to the outside was not a major impediment. As of the afternoon of August 14, the Goodman HTS was close to operational, but unfortunately enough ice remained on the dome that it could not be opened safely.

The last of the dome ice was removed August 15, all systems within the dome were operational, and observations were possible thereafter, though still without external network connections. Some observations were performed over the weekend (weather permitting).

On August 17, power for Pachón summit communications was restored and we were able to resume remote observing. As of August 18, operations were near-normal; we continue bringing instruments on-line based on next scheduled use, and should have all of them operational within the next few days. Carry-all service has resumed as far as the Pachón hotel.

No further updates will be posted.

We want to acknowledge the efforts of the NOAO-S road crew in re-opening the road under adverse conditions:


Aftermath of the storm: Left, snow outside old dining hall/clinic just below Gemini; this side of the
building is elevated about a meter off the ground. Right, road next to Gemini dome after clearing.
Also note the fog in both pictures.

Updated on August 19, 2021, 7:38 am